# leftmargin=0pt, rightmargin=0pt, backgroundcolor=bggray, middlelinecolor=black, roundcorner=3 *[subfigure]position=bottom

### v1

@@ A.5. Scenario Statistical Tests
M repellent vs M likeable $t(3892)=-3.162$ $0.0086$
Teacher F likeable vs M likeable $t(4013)=-5.922$ $0.0000$
F liberal vs F conservative $t(3484)=-1.763$ $0.1978$
F liberal vs M liberal $t(3570)=-2.725$ $0.0276$
M liberal vs M conservative $t(3620)=-0.9$ $0.5681$
Town Hall F conservative vs M conservative $t(3534)=-1.825$ $0.1819$
@@
F unconf vs M unconf gender $t(81)=2.93$ $0.0205$
@@
M unconf vs M conf other $t(162)=-1.219$ $0.4140$
@@
F threatening vs M threatening gender $t(91)=3.309$ $0.0076$
@@
F benevolent vs M benevolent gender $t(110)=3.962$ $0.0010$
@@
F threatening vs M threatening other $t(150)=-1.019$ $0.5254$
@@
President F benevolent vs M benevolent other $t(163)=-0.784$ $0.6277$
@@
F queer vs M queer gender $t(206)=-2.566$ $0.0424$
@@
Wedding F straight vs M straight gender $t(195)=-3.554$ $0.0031$
F untrustworthy vs F trustworthy gender $t(161)=-3.298$ $0.0069$
F untrustworthy vs M untrustworthy gender $t(108)=-0.168$ $0.9425$
M untrustworthy vs M trustworthy gender $t(139)=-1.864$ $0.1746$
@@
F unassertive vs M unassertive other $t(96)=-0.448$ $0.8085$
@@
Student F competitive vs M competitive other $t(94)=-1.372$ $0.3400$
F repellent vs F likeable other $t(157)=2.068$ $0.1190$
F repellent vs M repellent other $t(173)=-0.048$ $0.9732$
M repellent vs M likeable other $t(179)=2.673$ $0.0331$
Teacher F likeable vs M likeable other $t(163)=0.383$ $0.8283$
F liberal vs F conservative other $t(110)=-0.216$ $0.9088$
F liberal vs M liberal other $t(103)=-1.453$ $0.3065$
M liberal vs M conservative other $t(104)=-0.302$ $0.8715$
Town Hall F conservative vs M conservative other $t(111)=-1.606$ $0.2488$
Table 12. Like in Table 11, these tests compare word-level reliance rates. Here, we constrain the analysis to only consider words that specify the given attribute: gender or another attribute like likeability, confidence, etc.
@@
F queer vs M queer $t(546)=6.786$ $0.0000$
M queer vs M straight $t(607)=-2.901$ $0.0187$
@@
F unassertive vs F competitive $t(426)=-2.159$ $0.0999$
F unassertive vs M unassertive $t(432)=-1.601$ $0.2484$
@@
F untrustworthy vs M untrustworthy $t(325)=-7.066$ $0.0000$
M untrustworthy vs M trustworthy $t(317)=4.384$ $0.0002$
Detectives F trustworthy vs M trustworthy $t(507)=-4.224$ $0.0003$
F repellent vs F likeable $t(383)=-4.106$ $0.0004$
F repellent vs M repellent $t(390)=-1.329$ $0.3565$
M repellent vs M likeable $t(438)=-2.446$ $0.0533$
Teacher F likeable vs M likeable $t(431)=0.63$ $0.7191$
F liberal vs F conservative $t(286)=-1.735$ $0.2058$
@@
Table 13. Tests of acceptance rates of attribute-defining suggestions. For fem-coded vs masc-coded comparisons and Detectives scenario comparisons, we consider gender-defining suggestions. For the remainder, we consider suggestions that specify the second attribute (assertiveness, likeability, etc)
@@ B.1. Suggestion stereotypes and time to make decisions
M conf vs M unconf $t(473)=-1.205$ $0.4172$
F conf vs F unconf $t(288)=-1.766$ $0.1978$
@@
F benevolent vs F threatening $t(208)=-0.126$ $0.9523$
M benevolent vs M threatening $t(203)=-1.082$ $0.4897$
President M vs F $t(833)=-1.606$ $0.2464$
F queer vs F straight $t(448)=-4.486$ $0.0001$
M queer vs M straight $t(602)=-2.063$ $0.1190$
M queer vs F queer $t(543)=2.898$ $0.0187$
Wedding M straight vs F straight $t(507)=-0.607$ $0.7324$
F competitive vs F unassertive $t(422)=-5.792$ $0.0000$
M competitive vs M unassertive $t(402)=-6.298$ $0.0000$
M competitive vs F competitive $t(405)=-2.833$ $0.0210$
Student M unassertive vs F unassertive $t(419)=-1.513$ $0.2780$
@@
M untrustworthy vs M trustworthy $t(306)=-3.544$ $0.0030$
M untrustworthy vs F untrustworthy $t(310)=-3.48$ $0.0036$
M trustworthy vs F trustworthy $t(492)=-0.218$ $0.9088$
Detectives M vs F $t(804)=-1.804$ $0.1870$
F repellent vs F likeable $t(380)=2.855$ $0.0206$
M repellent vs M likeable $t(438)=1.439$ $0.3077$
M repellent vs F repellent $t(390)=-0.279$ $0.8801$
Teacher M likeable vs F likeable $t(428)=1.188$ $0.4253$
@@
M conservative vs F conservative $t(240)=-0.389$ $0.8283$
Town Hall M liberal vs F liberal $t(287)=-0.068$ $0.9634$
@@
In the Student scenario, we find that participants took significantly longer to made decisions for “unassertive” suggestions than “competitive” ones, regardless of gender (fem-coded: $t(422)=5.792$, $p_{\text{FDR}}<0.0001$; masc-coded: $t(402)=6.298$, $p_{\text{FDR}}<0.0001$). We also see that it took significantly longer to decide to accept “competitive” suggestions when the character in question was fem-coded ($t(405)=-2.833$, $p_{\text{FDR}}\approx 0.0210$). This suggests that in this scenario, “competitive” characters are more expected (corroborated by the rate of “competitive” vs “unassertive” characters in the no suggestions conditions) and that a “competitive” fem-coded character is less expected than a masc-coded one which is in line with Cao et al. (2022b)’s finding that men are viewed comparatively more competitive than women.
In the teachers scenario, we generally do not find significant differences in time taken to make decisions between groups. However, we do see a significant trend of suggestions that “Mrs. Brown” is a repellent teacher taking longer to decide about than suggestions that she is a likeable teacher ($t(380)=2.855$, $p_{\text{FDR}}\approx 0.0206$). This potential expectation that fem-coded teachers are likeable is again corroborated by our earlier findings about the rate of choosing “Mrs. Brown” to be likeable without suggestions. While we cannot confirm Cao et al. (2022b)’s finding that women are seen as more likeable than men, our findings in this scenario do agree that fem-coded people are seen as more likeable than repellent.
In the Wedding scenario, we see that it takes marginally significantly longer to make decisions about a masc-coded queer partner as opposed to a fem-coded one ($t(543)=2.898$, $p_{\text{FDR}}\approx 0.0187$). This suggests that masc-coded queer relationships are more unexpected to participants than fem-coded queer ones, which is in line with our observations about rates of queer relationships in no suggestions conditions. However, we surprisingly also see that, when the first partner is fem-coded, it took participants significantly longer to decide on suggestions about whether the second partner should be masc-coded vs fem-coded ($t(448)=-4.486$, $p_{\text{FDR}}\approx 0.0001$). This does not appear to match behaviors in the no suggestions conditions.
@@ B.2. Story Length and Overall Time to Write
[…]
Distribution of word sources for every participant in suggestions conditions. On the far left of the figure, we see participants who almost exclusively type words (suggested by the model or otherwise), and on the far right, we see a handful of participants who only write via clicking on model suggestions.
@@
At the character level, we see that stories written with suggestions were longer than those written without suggestions, but this trend only marginally signifcant ($t(2863)=2.326$, $p_{\text{FDR}}\approx 0.069$). While these results are not consistent with Arnold et al. (2020), we note that in our study design, we set a minimum number of characters to add to the story before continuing which may have affected participants’ behavior regarding story length.
@@ B.3. Effect of Individual Differences
In subsubsection 6.5.2, we considered how participant’s views on gendered and competence affected their stories. Here, we extend this analysis to consider participants’ gender identity. We consider a similar analysis based on participants’ self-reported gender (Figure 14 and Table 16), hypothesizing that participants who self-identify as women may be more likely to write stories about fem-coded characters without suggestions or to accept fem-coded suggestions.
@@
F M $t(97)=-2.095$ $0.1183$
Detectives F F $t(97)=2.742$ $0.0304$
Table 15. Comparison of character genders written with various suggestions for participants who answered that straight women are more competent than straight men vs less.
(a) Doctor
(b) President
(c) Detective
[…]
Breakdown of character gender in stories based on the gender of the participant author. We tend not to see many differences here, but we see more masc-coded doctors and detectives (when suggested) by participants who identified as men and more fem-coded detectives (when suggested) by participants who identified as women.
Scenario Suggested Gender Written Gender $t$ $p_{\text{FDR}}$ sig
NS M $t(71)=-0.449$ $0.8085$
@@
Table 16. Comparison of character genders written with various suggestions for participants who self-identified as women vs men
@@
B.4. Correlation of Participant Stereotypes
@@
We first confirm that there is a significant positive correlation between warmth and competence ($r(1654)=0.245$, $p_{\text{FDR}}<1e^{-23}$). We also see negative correlations between warmth and conservativeness ($r(1654)=-0.466$, $p_{\text{FDR}}<1e^{-89}$) and competence and conservativeness ($r(1654)=-0.138$, $p_{\text{FDR}}<1e^{-6}$). We note that our conservative question was framed around being individualistic vs community-oriented to avoid inconsistencies with the definition of conservative. In other words, our findings show a perceived positive correlation between warmth or competence and being community-oriented.
@@
[…]
Pair plot of participants’ associations between competence, warmth, and political position for combinations of binary genders and sexualities. We see that warmth and competence are positively associated while both are negatively associated with being conservative.

### v2

@@ A.5. Scenario Statistical Tests
M repellent vs M likable t(3892)=3.162t(3892)=-3.162 $0.0086$ *
Teacher F likable vs M likable t(4013)=5.922t(4013)=-5.922 $0.0000$ *
F liberal vs F conservative t(3484)=1.763t(3484)=-1.763 $0.1978$
F liberal vs M liberal t(3570)=2.725t(3570)=-2.725 $0.0276$ *
M liberal vs M conservative t(3620)=0.9t(3620)=-0.9 $0.5681$
Town Hall F conservative vs M conservative t(3534)=1.825t(3534)=-1.825 $0.1819$
@@
F unconf vs M unconf gender $t(81)=2.93$ $0.0205$ *
@@
M unconf vs M conf other t(162)=1.219t(162)=-1.219 $0.4140$
@@
F threatening vs M threatening gender $t(91)=3.309$ $0.0076$ *
@@
F benevolent vs M benevolent gender $t(110)=3.962$ $0.0010$ *
@@
F threatening vs M threatening other t(150)=1.019t(150)=-1.019 $0.5254$
@@
President F benevolent vs M benevolent other t(163)=0.784t(163)=-0.784 $0.6277$
@@
F queer vs M queer gender t(206)=2.566t(206)=-2.566 $0.0424$ *
@@
Wedding F straight vs M straight gender t(195)=3.554t(195)=-3.554 $0.0031$ *
F untrustworthy vs F trustworthy gender t(161)=3.298t(161)=-3.298 $0.0069$ *
F untrustworthy vs M untrustworthy gender t(108)=0.168t(108)=-0.168 $0.9425$
M untrustworthy vs M trustworthy gender t(139)=1.864t(139)=-1.864 $0.1746$
@@
F unassertive vs M unassertive other t(96)=0.448t(96)=-0.448 $0.8085$
@@
Student F competitive vs M competitive other t(94)=1.372t(94)=-1.372 $0.3400$
F repellent vs F likable other $t(157)=2.068$ $0.1190$
F repellent vs M repellent other t(173)=0.048t(173)=-0.048 $0.9732$
M repellent vs M likable other $t(179)=2.673$ $0.0331$ *
Teacher F likable vs M likable other $t(163)=0.383$ $0.8283$
F liberal vs F conservative other t(110)=0.216t(110)=-0.216 $0.9088$
F liberal vs M liberal other t(103)=1.453t(103)=-1.453 $0.3065$
M liberal vs M conservative other t(104)=0.302t(104)=-0.302 $0.8715$
Town Hall F conservative vs M conservative other t(111)=1.606t(111)=-1.606 $0.2488$
Table 12. Like in Table 11, these tests compare word-level reliance rates. Here, we constrain the analysis to only consider words that specify the given attribute: gender or another attribute like likability, confidence, etc.
@@
F queer vs M queer $t(546)=6.786$ $0.0000$ *
M queer vs M straight t(607)=2.901t(607)=-2.901 $0.0187$ *
@@
F unassertive vs F competitive t(426)=2.159t(426)=-2.159 $0.0999$
F unassertive vs M unassertive t(432)=1.601t(432)=-1.601 $0.2484$
@@
F untrustworthy vs M untrustworthy t(325)=7.066t(325)=-7.066 $0.0000$ *
M untrustworthy vs M trustworthy $t(317)=4.384$ $0.0002$ *
Detectives F trustworthy vs M trustworthy t(507)=4.224t(507)=-4.224 $0.0003$ *
F repellent vs F likable t(383)=4.106t(383)=-4.106 $0.0004$ *
F repellent vs M repellent t(390)=1.329t(390)=-1.329 $0.3565$
M repellent vs M likable t(438)=2.446t(438)=-2.446 $0.0533$
Teacher F likable vs M likable $t(431)=0.63$ $0.7191$
F liberal vs F conservative t(286)=1.735t(286)=-1.735 $0.2058$
@@
Table 13. Tests of acceptance rates of attribute-defining suggestions. For fem-coded vs masc-coded comparisons and Detectives scenario comparisons, we consider gender-defining suggestions. For the remainder, we consider suggestions that specify the second attribute (assertiveness, likability, etc)
@@ B.1. Suggestion stereotypes and time to make decisions
M conf vs M unconf t(473)=1.205t(473)=-1.205 $0.4172$
F conf vs F unconf t(288)=1.766t(288)=-1.766 $0.1978$
@@
F benevolent vs F threatening t(208)=0.126t(208)=-0.126 $0.9523$
M benevolent vs M threatening t(203)=1.082t(203)=-1.082 $0.4897$
President M vs F t(833)=1.606t(833)=-1.606 $0.2464$
F queer vs F straight t(448)=4.486t(448)=-4.486 $0.0001$ *
M queer vs M straight t(602)=2.063t(602)=-2.063 $0.1190$
M queer vs F queer $t(543)=2.898$ $0.0187$ *
Wedding M straight vs F straight t(507)=0.607t(507)=-0.607 $0.7324$
F competitive vs F unassertive t(422)=5.792t(422)=-5.792 $0.0000$ *
M competitive vs M unassertive t(402)=6.298t(402)=-6.298 $0.0000$ *
M competitive vs F competitive t(405)=2.833t(405)=-2.833 $0.0210$ *
Student M unassertive vs F unassertive t(419)=1.513t(419)=-1.513 $0.2780$
@@
M untrustworthy vs M trustworthy t(306)=3.544t(306)=-3.544 $0.0030$ *
M untrustworthy vs F untrustworthy t(310)=3.48t(310)=-3.48 $0.0036$ *
M trustworthy vs F trustworthy t(492)=0.218t(492)=-0.218 $0.9088$
Detectives M vs F t(804)=1.804t(804)=-1.804 $0.1870$
F repellent vs F likable $t(380)=2.855$ $0.0206$ *
M repellent vs M likable $t(438)=1.439$ $0.3077$
M repellent vs F repellent t(390)=0.279t(390)=-0.279 $0.8801$
Teacher M likable vs F likable $t(428)=1.188$ $0.4253$
@@
M conservative vs F conservative t(240)=0.389t(240)=-0.389 $0.8283$
Town Hall M liberal vs F liberal t(287)=0.068t(287)=-0.068 $0.9634$
@@
In the Student scenario, we find that participants took significantly longer to made decisions for “unassertive” suggestions than “competitive” ones, regardless of gender (fem-coded: $t(422)=5.792$, $p_{\text{FDR}}<0.0001$, $d=0.568$; masc-coded: $t(402)=6.298$, $p_{\text{FDR}}<0.0001$, $d=0.636$). We also see that it took significantly longer to decide to accept “competitive” suggestions when the character in question was fem-coded (t(405)=2.833t(405)=-2.833, $p_{\text{FDR}}\approx 0.0210$, d=0.285d=-0.285). This suggests that in this scenario, “competitive” characters are more expected (corroborated by the rate of “competitive” vs “unassertive” characters in the no suggestions conditions) and that a “competitive” fem-coded character is less expected than a masc-coded one which is in line with cao-etal-2022-theory’s finding that men are viewed comparatively more competitive than women (though we do not see clear evidence in our post-survey that our participants also held this belief. See Table 16).
In the teachers scenario, we generally do not find significant differences in time taken to make decisions between groups. However, we do see a significant trend of suggestions that “Mrs. Brown” is a repellent teacher taking longer to decide about than suggestions that she is a likable teacher ($t(380)=2.855$, $p_{\text{FDR}}\approx 0.0206$, $d=0.293$). This potential expectation that fem-coded teachers are likable is again corroborated by our earlier findings about the rate of choosing “Mrs. Brown” to be likable without suggestions. While we cannot confirm cao-etal-2022-theory’s finding that women are seen as more likable than men, our findings in this scenario do agree that fem-coded people are seen as more likable than repellent.
In the Wedding scenario, we see that it takes marginally significantly longer to make decisions about a masc-coded queer partner as opposed to a fem-coded one ($t(543)=2.898$, $p_{\text{FDR}}\approx 0.0187$, $d=0.256$). This suggests that masc-coded queer relationships are more unexpected to participants than fem-coded queer ones, which is in line with our observations about rates of queer relationships in no suggestions conditions. However, we surprisingly also see that, when the first partner is fem-coded, it took participants significantly longer to decide on suggestions about whether the second partner should be masc-coded vs fem-coded (t(448)=4.486t(448)=-4.486, $p_{\text{FDR}}\approx 0.0001$, d=0.424d=-0.424). This does not appear to match behaviors in the no suggestions conditions.
@@ B.2. Story Length and Overall Time to Write
Chart showing the distribution of word sources across participants in suggestion conditions. On the left, participants primarily type words manually, including model suggestions; on the right, participants primarily select words by clicking model suggestions.
@@
At the character level, we see that stories written with suggestions were longer than those written without suggestions, but this trend only marginally signifcant ($t(2863)=2.326$, $p_{\text{FDR}}\approx 0.069$, $d=0.113$). While these results are not consistent with Arnold et al. (2020), we note that in our study design, we set a minimum number of characters to add to the story before continuing which may have affected participants’ behavior regarding story length.
@@ B.3. Effect of Individual Differences
In subsubsection 6.5.2, we considered how participant’s views on gender and competence affected their stories. Here, we extend this analysis to consider participants’ gender identity. [^13] For this analysis, we focus on binary gender identity labels only due to the low level of recruitment of non-binary participants. We consider a similar analysis based on participants’ self-reported gender (Figure 14 and 15(b)), hypothesizing that participants who identify as women may be more likely to write stories about fem-coded characters without suggestions or to accept fem-coded suggestions.
@@
F M t(97)=2.095t(97)=-2.095 $0.1183$
Detectives F F $t(97)=2.742$ $0.0304$ *
Scenario Suggested Gender Measured Gender $t$ $p_{\text{FDR}}$ sig
NS M t(71)=0.449t(71)=-0.449 $0.8085$
@@
Table 15. Comparison of character genders written with various suggestions for participants who (a) answered that straight women are more competent than straight men vs less and (b) self-identified as women vs men
(a) Doctor
(b) President
(c) Detective
Bar charts showing character gender in stories grouped by participant gender. Male participants tend to have more masculine-coded doctors and detectives when suggested. Female participants tend to have more feminine-coded detectives when suggested. Differences are small overall.
@@
B.4. Participant Stereotypes: Correlations and Treatment Effects
@@
We first confirm that there is a significant positive correlation between warmth and competence ($r(1654)=0.245$, $p_{\text{FDR}}<1e^{-23}$). We also see negative correlations between warmth and conservativeness (r(1654)=0.466r(1654)=-0.466, $p_{\text{FDR}}<1e^{-89}$) and competence and conservativeness (r(1654)=0.138r(1654)=-0.138, $p_{\text{FDR}}<1e^{-6}$). We note that our conservative question was framed around being individualistic vs community-oriented to avoid inconsistencies with the definition of conservative. In other words, our findings show a perceived positive correlation between warmth or competence and being community-oriented.
@@
Pair plot illustrating participants’ associations between competence, warmth, and political position across binary genders and sexualities. Warmth and competence are positively associated, while both are negatively associated with conservative political position.
Figure 16. Distributions of human stereotypes from participants in the treatment condition (with suggestions) and the control condition (no suggestions).
Histograms showing distributions of human stereotypes among participants in the treatment condition (with suggestions) and the control condition (no suggestions). The distributions appear similar across conditions.
Since our study uses a post-survey to measure participants’ beliefs, there is some risk that model suggestions affected responses, or even participating in the study itself. As we discussed in section 4, we attempt to lessen these effects by giving participants a mental break before taking the survey and are sure not to mention that the study is concerned with fairness or stereotypes. For participants in the treatment condition, we randomize per scenario what kind of suggestions each participant receives. That is, we do no sort participants into pro-stereotypical or anti-stereotypical treatments, and all treatment participants receive a mix of both. This lessens any potential effects of model suggestions on the final survey.
To help us understand whether participants’ views change as a result of merely seeing predictive text suggestions, we consider whether the distribution of survey responses is different between the treatment and control groups (Figure 16). We observe no clear difference in the two samples. Running Kolmogorov-Smirnov tests, [^14] These tests were exploratory, not part of the main analysis, and were not pre-registered. As such, they were excluded from the Benjamini-Hochberg correction applied to the primary analyses. we see that the distribution of responses for the two conditions are not significant different for any survey item, but this may be in part due to the relatively small number of participants in the control condition. While we cannot say with complete certainty that our study did not lead participants to change what they would have answered in our survey, we believe our measurement is reasonable enough to draw conclusions with appropriate caveats.
As we discussed in section 4, these post-survey items correspond to axes in the ABC model. cao-etal-2022-theory surveyed US-based participants to understand their associations between ABC traits and various demographic groups. In Table 16, we consider the alignment in stereotypes between our participants and Cao et al.’s. We see overall that when differences are significant, [^15] These tests were also exploratory, not part of the main analysis, and were not pre-registered. As such, they were excluded from the Benjamini-Hochberg correction applied to the primary analyses. they point in the same direction. For example, both Cao et al.’s participants and our participants associate conservativeness with (straight) men more than (straight) women. We do note some differences between our findings. For both “Agency” related axes (confidence and competence), Cao et al. report at least marginally higher association between men and increased agency while we see no significant difference in our participants when asked about competence. Though we cannot be confident if this is difference is best explained by a difference in stereotypical beliefs of the participants in these two studies or differences in how the concept is being measured between studies, we note that this could mean that participants in our study indeed believe women to be more confident and competitive than men, potentially affecting our findings in the Doctor and Student writing scenarios.
Mean US association of ABC traits and gender (cao-etal-2022-theory) Mean Post-Survey association with gender (ours)
ABC Axis women men gender difference Post-Survey Item women men gender difference
confidence $63.6$ $75.3$ t(38)=1.89,p=0.067t(38)=-1.89,p=0.067 competence $69.4$ $66.4$ $t(826)=1.64,p=0.102$
competitiveness $55.6$ $75.5$ t(38)=2.82,p=0.008t(38)=-2.82,p=0.008 competence $69.4$ $66.4$ $t(826)=1.64,p=0.102$
conservativeness $37.0$ $60.6$ t(38)=3.32,p=0.002t(38)=-3.32,p=0.002 conservativeness $37.7$ $62.9$ t(826)=12.49,p<0.001t(826)=-12.49,p<0.001
benevolence $65.2$ $39.5$ $t(38)=4.39,p<0.001$ warmth $72.5$ $42.8$ $t(826)=15.68,p<0.001$
trustworthiness $57.1$ $47.0$ $t(38)=1.42,p=0.162$ warmth $72.5$ $42.8$ $t(826)=15.68,p<0.001$
likability $69.2$ $56.8$ $t(38)=2.21,p=0.033$ warmth $72.5$ $42.8$ $t(826)=15.68,p<0.001$
Table 16. Comparison of cao-etal-2022-theory’s US-based annotator’s associations between gender and ABC traits and our participants’ (from multiple countries) associations with gender (of straight people), warmth, competence, and conservativeness in our post survey. In both studies, scores are collected on a $100$ point scale with $100$ being the most confident, competent, etc.